The ratio of the number of sides of three regular polygons is 4:5:6. What is the ratio of the measures of their exterior angles?

Answer by stanbon(75887)

The ratio of the number of sides of three regular polygons is 4:5:6. What is the ratio of the measures of their exterior angles?
-----
4 sides: each ext. angle = 360/4 = 90 degrees
5 sides: each ext. angle = 360/5 = 72 degrees
6 sides: each ext. angle = 360/6 = 60 degrees
------
Answer: 90:72:60
===================
Cheers,
Stan H.
===================